# MetaMixer Is All You Need

### v1

@@ Appendix D Effective Receptive Field (ERF)
Previous works [24, 54] have identified the Effective Receptive Field (ERF) as a key factor in enhancing CNN performance, scaling kernels up to 31$\times$31 or even 51$\times$51. In this section, we compare the ERFs of ConvNeXt and Swin Transformer, which use the same kernel (window) size, with our FFNet. Following RepLKNet [24], we sample and resize 50 images from the ImageNet validation set to 1024 $\times$ 1024, and measure the contribution of pixels on the input images to the central point of the feature map produced by the last layer. Contributions are aggregated and mapped onto a 1024 $\times$ 1024 matrix. We further quantify the ERF of each model by calculating the high-contribution area ratio $r$ of a minimum rectangle that encompasses the contribution scores over a certain threshold $t$. As depicted in Fig. 9, while high-contribution pixels in ConvNeXt and Swin are concentrated around the center of the input, FFNet’s high-contribution pixels are dispersed across a wider ERF. These results demonstrate that our model more effectively utilizes convolutions with the same kernel size to consider a larger range of pixels for final predictions. These results are consistent with the examples in Fig. 4.
Table 8: Robustness evaluation of FFNet. “*” indicates that model is trained using DeiT [89] distillation.
Model #Params. GPU Throughput Mobile Latency IN-Clean IN-A [37] IN-R [35] IN-SK [96] IN-V2 [75] (M) (img/sec) (ms) Top-1 (%) Top-1 (%) Top-1 (%) Top-1 (%) Top-1 (%) FastViT-SA12 [93] 10.9 731 2.0 80.6 17.2 42.6 29.7 - FFNet-1 13.7 1113 1.8 81.3 18.8 44.1 31.6 70.1 FFNet-1 13.7 1113 1.8 82.1 24.2 47.0 33.5 71.5 Swin-T [59] 29.0 - 24.8 81.3 21.6 41.3 29.1 69.7 ConvNeXt-T [60] 28.6 789 3.7 82.1 24.2 47.2 33.8 71.0 FocalNet-T (LRF) [108] 28.6 599 4.3 82.3 23.5 45.1 31.8 71.2 EfficientNet-B4 [85] 19.3 572 5.2 82.9 26.3 47.1 34.1 72.3 FFNet-2 26.9 674 3.1 82.9 27.3 48.0 34.6 71.8 FFNet-2 26.9 674 3.1 83.7 33.4 51.3 37.8 73.3 Swin-B [59] 87.8 - 33.5 83.5 35.8 46.6 32.4 72.3 ConvNeXt-S [60] 50.2 459 5.0 83.1 31.3 49.6 37.1 72.5 ConvNeXt-B [60] 88.6 399 8.1 83.8 36.7 51.3 38.2 73.7 FastViT-MA36 [93] 42.7 314 4.7 83.9 34.6 49.5 36.6 - FocalNet-B (LRF) [108] 88.7 278 10.0 83.9 38.3 48.1 35.7 73.5 FFNet-3 48.3 490 4.5 83.9 36.2 50.8 37.6 73.4 FFNet-3 48.3 490 4.5 84.5 39.5 54.1 40.2 74.5 ImageNet-1K 384 $\times$ 384 fine-tuned models Swin-B [59] 87.8 - 181.6 84.5 42.0 47.2 33.4 73.2 ConvNeXt-B [60] 88.6 220 21.3 85.1 45.6 52.9 39.5 75.2 FFNet-4 79.2 264 15.2 85.3 52.3 54.8 41.1 76.0
Appendix E Robustness Evaluation
In this section, we test our models on several robustness benchmark datasets: ImageNet-A [37], a test set comprising natural adversarial examples; ImageNet-R [35], an extended test set containing samples that are misclassified by ResNet-50 [34]; ImageNet-Sketch [96], comprised solely of hand-drawn images, serves as an effective test set for classification capabilities in the absence of texture or color information; and ImageNet-V2 [96], an extended test set that employs the same sampling strategy as ImageNet-1K.
Robustness evaluation results are presented in Tab. 8. Our models exhibit promising robust performance, outperforming well-tuned models on four additional test sets. With distillation, FFNets show strong generalization capabilities across all scales. Notably, on the most challenging test set, ImageNet-A, FFNet-4 achieves a 52.3% accuracy, surpassing Swin-B and ConvNeXt-B by 10.3% and 6.7%, respectively.
Appendix F Structural Re-parameterization
Figure 10: Visual explanation of re-parameterizing multi-branches for extracting multi-scale features. For a detailed discussion, refer to Appendix F.
Structural re-parameterization [25] is a methodology that transforms train-time multi-branch architectures into plain architectures through parameter transformation. We utilize this methodology to incorporate small kernels (e.g., 3$\times$3 or 5$\times$5) into larger ones (e.g., 7$\times$7 or 9$\times$9), as shown in Fig. 10. Additionally, for semantic segmentation model FFNetseg, we add 9$\times$1 and 1$\times$9 convolution branches to effectively extract features of strip-like objects, such as human and tree. This approach enables our mixer to capture patterns of various scales and aspect ratios.
Using multi-branches can be interpreted as one-to-many interactions, similar to the self-attention, leveraging multiple key-value pairs for a single query. However, unlike self-attention, which typically considers a large number of key-values (e.g., 196 or 49 tokens) per query, our method uses only 2 or 4 key-values, making element-wise addition followed by GELU (“collaborative weighting”) more natural than softmax (“competitive weighting”). Meanwhile, based on FFNified attention, leveraging softmax gating with multi-branches for context-dependent key-values [9] or handling multiple tasks [122] presents an intriguing direction for future research. As demonstrated in Tab. 9, employing the re-parameterization technique within our mixer design effectively boosts performance without additional inference costs. However, this technique results in increased training time due to the added branches.
Variant Ablation Train Time (hrs) Top-1 Acc. (%) FFNet-1 plain architecture 34.5 81.0 multi-branches w/ re-param. 36.6 81.3 FFNet-2 plain architecture 73.2 82.5 multi-branches w/ re-param. 77.5 82.9
Appendix G Architecture Details
G.1 ImageNet Classification Model Architecture Details
@@
The detailed architecture hyper-parameters are shown in Tab. 10. Given an input image, we first apply two 3$\times$3 strided convolutions to it. Batch normalization and GELU are used after each 3$\times$3 convolution in stem. Then, the tokens pass through four stages of stacked FFNet blocks for hierarchical feature extraction. In the early stages, 3$\times$3 convolutions are primarily used to capture high-frequency local information, followed predominantly by 7$\times$7 convolutions in later stages for capturing low-frequency global information [81]. When employing 7$\times$7 or 9$\times$9 convolutions, we leverage the train-time multi-branches with re-parameterization technique introduced in Appendix F. The kernel size for depthwise convolution in the channel mixer is set to 7$\times$7 for all models except FFNet-1 and $\text{FFNet}_{\textit{seg}}$, which use 3$\times$3.
G.2 Super-Resolution Model Architecture Details
@@
In this section, we explain the architecture of the proposed $\text{FFNet}{{}_{\textit{sr}}}$, adapted for the Super-Resolution (SR) task, as shown in Fig. 11. $\text{FFNet}{{}_{\textit{sr}}}$ consists of three main components: the Shallow Feature Extractor ($H_{SF}$), the Deep Feature Extractor ($H_{DF}$), and the Upscaling Module ($H_{Rec}$).
@@
We utilize an anchor-based residual skip for the long residual skip, where the input LR image is repeated 4 times along RGB channel-wise [28, 125]. After passing through $H_{Rec}$, the repeated LR image becomes equivalent to a nearest-neighbor interpolated image, enabling the feature extractors to concentrate on recovering high-frequency details.
The upscaling module, $H_{Rec}$, takes the deep features and the long-residual skip, adds them together, and then upscales the result to reconstruct the High-Resolution image. It uses a single PixelShuffle [80] operation, which rearranges the channel dimensions into spatial dimensions in proportion to the upscaling factor. This approach is more efficient and effective than the Transposed Convolution, which often results in checkerboard artifacts [99].
G.3 3D Semantic Segmentation Model Architecture Details
We adopt the serialization-based point cloud preprocessing introduced in PTv3[103] and use a submanifold convolution with 5$\times$5$\times$5 kernel size for embeddings. The encoder and decoder depths are set at [2, 2, 6, 2] and [1, 1, 1, 1], respectively, with channel numbers of [64, 128, 256, 512] for the encoder and [64, 64, 128, 256] for the decoder. The expansion ratio for channel mixer is uniformly set at 4 across all blocks, with GELU as the activation function. Furthermore, we incorporate batch normalization after each submanifold convolution, in addition to the pre-norm, to further stabilize the variance of the feature maps.
G.4 Time Series Forecaster Architecture Details
We introduce two FFN variants tailored for time series data, which maintain the structural integrity of our FFNet block design (Fig. 3 (c)) while effectively computing cross-time and cross-variable interactions. Given the limitations of vanilla FFNs that operate solely along the channel axis, in a multivariate setting with both temporal and variable axes, we propose CVIFFN (Cross-Variable Interaction focused FFN) and CIFFN (Channel Interaction focused FFN) to extract cross-variable dependencies and channel information, respectively. Details of these modules, including PyTorch-like code and intuitive visual explanations, can be found in Algorithm 1, 2, and Fig. 12. Both CVIFFN and CIFFN capture cross-variable and channel information through the first Fully-Connected (FC) layer respectively, then aggregate features across both axes using appropriate group&\&&reshape operations and second FC layer. Thus, our FFNet block for time series replaces the basic block’s FFN with CIFFN and the token mixer’s 1$\times$1 convolution with CVIFFN.
By default, our model includes 1 FFNet block with a hidden dimension of 64 and an expansion ratio of 12. For patch embedding, the patch size and stride are set at 4 and 2, respectively. For the Exchange, ETTh1, and ETTh2 datasets, we reduce the expansion ratio to 2, 2, and 8, respectively, to prevent overfitting. To address distribution shift problem, we utilize RevIN [42]. Finally, for features $O\in\mathbb{R}^{M\times D\times N}$ , we flatten them along the channel $D$ and token $N$ axes, and then compute the final prediction using a linear layer with output channels equal to the prediction length.
@@ G.4 Time Series Forecaster Architecture Details
Appendix H Training Settings
H.1 3D Semantic Segmentation
We train all models using a grid size of 0.02m, a batch size of 12, and a drop path rate of 0.3. We employ the AdamW optimizer with an initial learning rate of $5\times 10^{-3}$, adjusted via cosine decay, and a weight decay of 0.05. Models are trained for 800 epochs on ScanNet series [17, 77] and 3000 epochs on S3DIS [1] datasets. The data augmentation strategies employed are identical to those used in the PT series [104, 103].
H.2 Long-Term Time Series Forecasting
We train our model with the L2 loss, using Adam optimizer with an initial learning rate of $10^{-4}$. The standard training procedure is 100 epochs, incorporating proper early stopping. Following [57], we fix lookback length 96 for all the datasets and all experiments are repeated three times. We also borrow from the baseline results from iTransformer [57]. Detailed dataset descriptions are summarized in Tab. 11.
@@ H.2 Long-Term Time Series Forecasting
Figure 13: Visual comparison for $\times$2 SR models. The patches for comparison are marked with red boxes. Zoom-in for better visibility.
Appendix I Super-Resolution Visual Results
In this section, we present visual comparisons of our $\text{FFNet}_{\textit{sr}}$ model against state-of-the-art SR models at $\times$2 scale. Figure 13 zooms in on each SR result to emphasize the differences in image sharpness and fine-grained detail preservation, providing corresponding PSNR values. These visualizations underscore the exceptional performance of $\text{FFNet}_{\textit{sr}}$ in the SR task, demonstrating task-generality of the MetaMixer framework.
Appendix J Limitations
While our proposed MetaMixer-based convolutional mixer demonstrates remarkable performance across multiple tasks, showcasing that the MetaMixer itself is crucial rather than specific modules like attention, its effectiveness in more diverse scenarios (such as large-scale datasets like ImageNet-21K or generative modeling requiring flexible conditional operations) remains unproven. Moreover, we acknowledge that our model’s performance is partially improved with the help of recent methods. However, we hope that the new perspectives and possibilities we present will serve as a starting point for future mixer design space and strengthen the systematic approach to model development.

### v2

@@ Appendix D Effective Receptive Field Analysis
Previous works [25, 55] have identified the Effective Receptive Field (ERF) as a key factor in enhancing CNN performance. In this section, we mainly compare the ERFs of ConvNeXt and Swin Transformer, which use the same kernel (window) size, with our FFNet. Following RepLKNet [25], we sample and resize 50 images from the ImageNet validation set to 1024 $\times$ 1024, and measure the contribution of pixels on the input images to the central point of the feature map produced by the last layer. Contributions are aggregated and mapped onto a 1024 $\times$ 1024 matrix. We further quantify the ERF of each model by calculating the high-contribution area ratio $r$ of a minimum rectangle that encompasses the contribution scores over a certain threshold $t$.
As depicted in Fig. 10, while high-contribution pixels in ConvNeXt and Swin are concentrated around the center of the input, FFNet’s high-contribution pixels are dispersed across a wider ERF. These results demonstrate that our model more effectively utilizes convolutions with the same kernel size to consider a larger range of pixels for final predictions. Moreover, FFNified attention exhibits a receptive field comparable to that of FastViT [94] and DeiT [90], which employ global self-attention. These results indicate that despite using only convolution, FFNified attention harnesses large kernels effectively to capture global context. These results are consistent with the examples in Fig. 4.
@@
Appendix E Structural Re-parameterization
Structural re-parameterization [24] is a methodology that transforms train-time multi-branch architectures into plain architectures through parameter transformation (Addition). We utilize this methodology to incorporate small kernels (e.g., 3$\times$3 or 5$\times$5) into larger ones (e.g., 7$\times$7 or 9$\times$9), as shown in Fig. 9. Additionally, for semantic segmentation model FFNetseg, we add 9$\times$1 and 1$\times$9 convolution branches to effectively extract features of strip-like objects, such as human and tree. This approach enables our mixer to capture patterns of various scales and aspect ratios without adding inference latency.
Interestingly, using multi-branches can be interpreted as one-to-many interactions, similar to the self-attention, leveraging multiple key-value pairs (kernels) for a single query (channel). However, unlike self-attention, which typically considers a large number of key-values (e.g., 196 or 49 tokens) per query, our method uses only 2 or 4 key-values, making element-wise addition followed by GELU (“collaborative weighting”) more natural than softmax (“competitive weighting”). Meanwhile, based on FFNified attention, leveraging softmax gating with multi-branches for context-dependent key-values [10] or handling multiple tasks [121] presents an intriguing direction for future work. As demonstrated in Tab. 10, employing the re-parameterization technique within our mixer design effectively boosts performance without additional inference costs. However, this technique results in increased training time due to the added branches.
Variant Ablation Train Time (hrs) Top-1 Acc. (%)
FFNet-1 plain architecture 34.5 81.0
multi-branches w/ re-param. 36.6 81.3
FFNet-2 plain architecture 73.2 82.5
multi-branches w/ re-param. 77.5 82.9
Appendix F Implementation details
F.1 Image Recognition
Architecture. The detailed architecture hyperparameters are shown in Tab. 9. Given an input image, we first apply two 3$\times$3 strided convolutions to it. Batch normalization and GELU are used after each 3$\times$3 convolution in stem. Then, the tokens pass through four stages of stacked FFNet blocks for hierarchical feature extraction. In the early stages, 3$\times$3 convolutions are primarily used to capture high-frequency local information, followed predominantly by 7$\times$7 convolutions in later stages for capturing low-frequency global information [81]. For downsampling, we utilize a 7$\times$7 strided depthwise convolution followed by a 1$\times$1 convolution. We also use LayerScale [91] to make our deep models more stable during training.
For system-level segmentation comparison, we introduce $\text{FFNet}_{\textit{seg}}$, a model that exclusively utilizes FFNet blocks across both the backbone and head. Recent works [5, 34, 116] demonstrate that long-range interactions and multi-scale features are crucial for competitive performance in segmentation models. Accordingly, to enlarge the receptive field, we employ 9$\times$9 kernels in the final two stages of the backbone and the head. Additionally, to effectively process features of objects with varying sizes, we combine small-kernel and strip convolutions [39, 71] (9$\times$1, 1$\times$9) with large-kernel convolution, enhancing flexible feature extraction. For this, when employing 7$\times$7 or 9$\times$9 convolutions, we leverage re-parameterization technique with the train-time multi-branches introduced in Appendix E. The kernel size for depthwise convolution in the channel mixer is set to 7$\times$7 for all models except FFNet-1 and $\text{FFNet}_{\textit{seg}}$, which use 3$\times$3 kernels.
@@ F.1 Image Recognition
F.2 Super-Resolution
Architecture. In this section, we explain the architecture of the proposed $\text{FFNet}{{}_{\textit{sr}}}$, adapted for the Super-Resolution (SR) task, as shown in Fig. 11. $\text{FFNet}{{}_{\textit{sr}}}$ consists of three main components: the Shallow Feature Extractor ($H_{SF}$), the Deep Feature Extractor ($H_{DF}$), and the Upscaling Module ($H_{Rec}$).
@@ F.2 Super-Resolution
We utilize an anchor-based residual skip for the long residual skip, where the input LR image is repeated $r^{2}$ times along RGB channel-wise [28, 124], where $r$ denotes the upscaling factor. After passing through $H_{Rec}$, the repeated LR image becomes equivalent to a nearest-neighbor interpolated image, enabling the feature extractors to concentrate on recovering high-frequency details.
The upscaling module, $H_{Rec}$, takes the deep features and the long-residual skip, adds them together, and then upscales the result to reconstruct the High-Resolution (HR) image. It uses a single PixelShuffle [80] operation, which rearranges the channel dimensions into spatial dimensions in proportion to the $r$. This approach is more efficient and effective than the Transposed Convolution, which often results in checkerboard artifacts [99].
Settings. In each training batch, we randomly crop LR patches of size 64$\times$64 as input and augment the patches with random horizontal flips and rotations. We use AdamW optimizer for training with a batch size of 64 for 1M iterations as follows previous study [118]. The initial learning rate is set to 0.001 and updated by the cosine annealing scheme. All latency and memory usage are measured corresponding to an HR image of size 1280$\times$720.
F.3 3D Semantic Segmentation
Architecture. We adopt the serialization-based point cloud preprocessing introduced in PTv3 [103] and use a submanifold convolution with 5$\times$5$\times$5 kernel size for embeddings. The encoder and decoder depths are set at [2, 2, 6, 2] and [1, 1, 1, 1], respectively, with channel numbers of [64, 128, 256, 512] for the encoder and [64, 64, 128, 256] for the decoder. The expansion ratio for channel mixer is uniformly set at 4 across all blocks, with GELU as the activation function. Furthermore, we incorporate batch normalization after each submanifold convolution, in addition to the pre-norm, to further stabilize the variance of the feature maps.
Settings. We train all models using a grid size of 0.02m, a batch size of 12, and a drop path rate of 0.3. We employ the AdamW optimizer with an initial learning rate of $5\times 10^{-3}$, adjusted via cosine decay, and a weight decay of 0.05. Models are trained for 800 epochs on ScanNet series [18, 77] and 3000 epochs on S3DIS [1] datasets. The data augmentation strategies employed are identical to those used in the PT series [102, 103].
@@
F.4 Time Series Forecasting
Architecture. We introduce two FFN variants tailored for time series data, which maintain the structural integrity of our FFNet block design (Fig. 3 (c)) while effectively computing cross-time and cross-variable interactions. Given the limitations of vanilla FFNs that operate solely along the channel axis, in a multivariate setting with both temporal and variable axes, we propose CVIFFN (Cross-Variable Interaction focused FFN) and CIFFN (Channel Interaction focused FFN) to capture cross-variable dependencies and channel information, respectively. Details of these modules, including PyTorch-like code and intuitive visual explanations, can be found in Algorithm 1, 2, and Fig. 12. Both CVIFFN and CIFFN capture cross-variable and channel information through the first Fully-Connected (FC) layer respectively, then aggregate features across both axes using appropriate group&\&&reshape operations and a second FC layer. Thus, our FFNet block for time series replaces the basic block’s FFN with CIFFN and the token mixer’s 1$\times$1 convolution with CVIFFN (see Fig. 12 (c)).
@@ F.4 Time Series Forecasting
By default, our model includes 1 FFNet block with a hidden dimension of 64 and an expansion ratio of 12. For patch embedding, the patch size and stride are set at 4 and 2, respectively. For the Exchange, ETTh1, and ETTh2 datasets, we reduce the expansion ratio to 1, 2, and 8, respectively, to mitigate overfitting. To address distribution shift problem, we also utilize RevIN [43]. Finally, for features $O\in\mathbb{R}^{M\times D\times N}$ , we flatten them along the channel $D$ and token $N$ axes, and then compute the final prediction using a linear layer with output channels equal to the prediction length.
Settings. We train our model with the L2 loss, using Adam optimizer with an initial learning rate of $10^{-4}$. The standard training procedure is 100 epochs, incorporating proper early stopping. Following [57], we fix lookback length 96 for all the datasets and all experiments are repeated three times. We also borrow from the baseline results from iTransformer [57]. Detailed dataset descriptions are summarized in Tab. 11.
Appendix G Qualitative Results
G.1 Super-Resolution
In this section, we present visual comparisons of our $\text{FFNet}_{\textit{sr}}$ model against state-of-the-art SR models at $\times$4 scale. Figure 13 zooms in on each SR result to emphasize the differences in image sharpness and fine-grained detail preservation, providing corresponding PSNR values. These visualizations underscore the exceptional performance of $\text{FFNet}_{\textit{sr}}$ in the SR task, demonstrating task-generality of the MetaMixer-based convolutional framework.
G.2 ImageNet Generation
We provide the generated samples of latent diffusion model with our hybrid mixer design in Fig. 14. Our results confirm that our proposed mixer design is also effective in generative modeling.
Figure 13: Visual comparison for $\times$4 SR models. The patches for comparison are marked with red boxes. Zoom-in for better visibility.
Figure 14: Selected samples generated using our U-FFNet ImageNet-256 model with classifier-free guidance.
Appendix H Limitations
While our proposed MetaMixer-based convolutional mixer demonstrates remarkable performance across multiple tasks, showcasing that the MetaMixer itself is crucial rather than specific modules like attention, its effectiveness in more diverse scenarios, such as large-scale datasets like ImageNet-21k, remains unproven. Moreover, we acknowledge that the performance of our model is partially improved with the help of previous techniques. However, we hope that our unified perspectives and possibilities we present will serve as a starting point for future mixer design space and strengthen the systematic approach to model development.